OpenStudy (anonymous):

I need the answer to this fast. Thanks! Wendy took a trip from city A to city B, a distance of 380 mi. She traveled part of the way by bus, which arrived at the train station just in time for Wendy to complete her journey by train. The bus averaged 40 mi/h, and the train averaged 120 mi/h. The entire trip took 7 1/2h. How long did Wendy spend on the train?

OpenStudy (anonymous):

Let x be the distance by bus and 380-x the distance by train, then \[ \frac x{40} + \frac{380-x}{120}= \frac {15}2 \\ \] Solve for x.

OpenStudy (anonymous):

I got 260 and thats incorrect

OpenStudy (anonymous):

\[ \frac x{40} + \frac{380-x}{120}= \frac {15}2 \\ 3 x + 380-x= 900\\ 2 x = 900-380\\ 2 x = 520\\ x= 260 \]

OpenStudy (anonymous):

Yes, you are right. How many miles per train?

OpenStudy (anonymous):

It says it is incorrect

OpenStudy (anonymous):

We are not done yet. How many miles per train? 380-260=120 miles by train. What is the speed of the train? You can now deduce the time spent on the train.

OpenStudy (anonymous):

The speed of the train is 120 mph. What is the time spent on the train to travel 120 miles?

OpenStudy (anonymous):

i hour?

OpenStudy (anonymous):

You are done.

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

yw